40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 623     Entry time: Wed Jul 2 13:56:10 2008
Author: Rob, Yoichi, John 
Type: Update 
Category: Locking 
Subject: 24.5 Hz resonance 
Work continues on trying to reduce the CARM offset using dc signals from PO_DC. Got up to arm powers of
~35 last night.

We found that progress was stymied by an oscillation around 24 Hz. This oscillation was clearly visible
in the intensity of the light at REFL, PO and TrX.

Initially we suspected that this oscillation was due to an instability in the CARM loop. We attempted to
solve the problem by tuning the crossover frequncy of the AO and MC_L paths and shaping the MC_L loop to
reduce the impact of the 24 Hz noise.

After some quick tests we found that the 24 Hz signal was present even when dc CARM was used. It appears
that the peak is in fact due to a SOS mechanical resonance. We currently suspect a roll mode.

We're going to check that PRC, MICH and DARM have filters to attenuate the 24 Hz line. We'll also look at the
SUS_POS bandstop filters to see where they are centred.

The ISS was behaving strangely again. Constantly saturated at 5dB of gain. Someone needs to look a this.
Attachment 1: locking080702.png  71 kB  | Hide | Hide all
locking080702.png
ELOG V3.1.3-